A person drives to a tennis match at 48mph and returns home at 54mph. If the total driving time was 17 hours, how far away is the match?
Answer by oberobic(2304)

x = driving time to the match
y = driving time from the match
x + y = 17
So x = 17-y and y = 17-x
Using the distance formula d = rt, we know the rates, but not the times or distances.
But we know the distance is same both ways (or at least we can assume it is).
Substituting what we know about the time...
48x = d
54(17-x) = d
So,
48x = 54(17-x)
divide both sides be 6
8x = 9(17-x)
expand
8x = 153 - 9x
add 9x to both sides
17x = 153
divide through by 17
x = 153/17 = 9
x = 9, which is the time driving to the match
y = 17-9 = 8, which is the time driving home
.
How far away is the match?
Well, we can use that to check our answer.
d = 48 miles/hr * 9 hr = 432 miles
d = 54 miles/hr * 8 hr = 432 miles
Correct!
.
Answer:
The match is 432 miles away.
.
Done.